You are considering taking the following option positions. As part of your analysis calculate the stock price or prices on expiration above which or below which the positions will be profitable (ignore dividends and interest). The current stock price is $68.00

(a) Buy a straddle with an exercise price of 70, where each option costs $5.00.

(b) Sell a straddle with an exercise price of 70, where each option costs $5.00.
Answer #1

1.
Lower price=70-5-5=60

Upper price=70+5+5=80

Price should be less than 60 or more than 80

2.
Price should be between 60 and 80